OpenStudy (anonymous):

help me to this one pleaseee. differentiation of algebraic function. y= 2x^2/ squareroot of x - cube root of x^2

OpenStudy (anonymous):

@matricked

OpenStudy (anonymous):

can you post the question with proper braces (if) otherwise answer differs

OpenStudy (anonymous):



OpenStudy (anonymous):

@matricked

OpenStudy (anonymous):

is your question (2x^2/sqrt(x)) - cbrt(x^2) or 2x^2 / ( sqrt(x) - cbrt(x^2) )

OpenStudy (anonymous):

the first one you wrote

OpenStudy (anonymous):

so the question can be re-written as y= 2*x^(3/2) - x^(2/3)

OpenStudy (anonymous):

yes. then whats should i do next?

OpenStudy (anonymous):

now differentiate wrt x

OpenStudy (anonymous):

dy/dx = 2*(3/2) *x^(3/2-1) - (2/3) *x^(2/3-1) dy/dx = 3*x^(1/2) - (2/3) * x^(-1/3)

OpenStudy (anonymous):

okay i understand now. should i substitute the whole number now

OpenStudy (anonymous):

is that the final answer? @matricked

OpenStudy (anonymous):

thankyou godbless

OpenStudy (anonymous):

yw